A real estate firm is building a new apartment tower (similar to its other developments). It will require spending $17 million today. This project will then generate $2 million in free cash flow each year for the next 20 years (from t=1 to t=20). The firm is financed half with equity and half with debt (it has no excess cash). The firm's equity beta is 1.2, debt yield to maturity is 6.0%, and tax rate is 20%. The expected return on the S&P500 Index is expected to be 7.0% each year and the relevant risk-free rate is 2.0%. What is the NPV of this project (in millions)?

A.-3.4

B.-1.9

C.0

D.1.1

E.3.0

F.5.2
Answer #1
1] After tax cost of debt = 6%*(1-20%) = 4.80%
Cost of equity per CAPM = 2%+1.2*(7%-2%) = 8.00%
WACC = 4.80%*0.5+8%*0.5 = 6.40%
2] NPV = 2*(1.064^20-1)/(0.064*1.064^20)-17 = $                     5.2
Answer: Option [F] $5.2 million